OpenStudy (df001):

How to find point (x,y) on a Unit Circle based in the reference angle of 5pi/4?

OpenStudy (df001):

The reference angle of \[\frac{ 5\pi }{ 4 } is \frac{ \pi }{ 4 }\]

OpenStudy (df001):

Hello

OpenStudy (df001):